# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

### v2

@@ Abstract
Adaptability is critical in cooperative multi-agent reinforcement learning (MARL), where agents must learn specialised or homogeneous behaviours for diverse tasks. While parameter sharing methods are sample-efficient, they often encounter gradient interference among agents, limiting their behavioural diversity. Conversely, non-parameter sharing approaches enable specialisation, but are computationally demanding and sample-inefficient. To address these issues, we propose HyperMARL, a parameter sharing approach that uses hypernetworks to dynamically generate agent-specific actor and critic parameters, without altering the learning objective or requiring preset diversity levels. By decoupling observation- and agent-conditioned gradients, HyperMARL empirically reduces policy gradient variance and facilitates specialisation within FuPS, suggesting it can mitigate cross-agent interference. Across multiple MARL benchmarks involving up to twenty agents—and requiring homogeneous, heterogeneous, or mixed behaviours—HyperMARL consistently performs competitively with fully shared, non-parameter-sharing, and diversity-promoting baselines, all while preserving a behavioural diversity level comparable to non-parameter sharing. These findings establish hypernetworks as a versatile approach for MARL across diverse environments. [^1] All code will be made available soon.
Multi-Agent Reinforcement Learning, Parameter Sharing, Specialisation, Behavioural Diversity
@@ 1 Introduction
Specialist and generalist behaviours are critical to collective intelligence. In biological and social systems, individual specialisation enhances performance by allowing agents to excel at specific tasks (Woolley et al., 2015; Smith et al., 2008; Surowiecki, 2004), while generalist (homogeneous) behaviours promote adaptability and efficiency (Kassen, 2002; Williams & O’Reilly III, 1998). In Multi-Agent Reinforcement Learning (MARL), tasks similarly vary and require adaptability (Li et al., 2021; Bettini et al., 2024; Albrecht et al., 2024): some demand specialised agent behaviours, others benefit from homogeneous strategies, and some may require a blend of both (mixed behaviours).
@@
Thus optimal performance is tied to being able to represent the required behaviours. No parameter sharing (NoPS) (Lowe et al., 2017) enables specialisation by using separate policy and critic networks per agent, but introduces significant computational overhead and sample inefficiency (Christianos et al., 2021). Conversely, full parameter sharing (FuPS) (Tan, 1993; Gupta et al., 2017; Foerster et al., 2016) tackles these inefficiencies by training a single shared network, but struggles in scenarios requiring diverse or specialised behaviours (Kim & Sung, 2023; Fu et al., 2022; Li et al., 2021).
@@
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards (Li et al., 2021; Jiang & Lu, 2021), role-based allocations (Wang et al., 2020a, b), specialised architectures (Kim & Sung, 2023; Li et al., 2024; Bettini et al., 2024), or sequential updates (Zhong et al., 2024). However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
@@
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without modifying the learning objective or relying on prior knowledge of the optimal diversity level? A key difficulty with FuPS (particularly for diverse behaviours) is the gradient interference among agents, whereby their updates negatively impact each other’s learning (Christianos et al., 2021; Zhong et al., 2024).
@@
To address this, we propose HyperMARL, a novel parameter sharing method that leverages hypernetworks (Ha et al., 2016) to generate agent-specific actor and critic parameters conditioned on agent IDs or embeddings (Figure 1). Hypernetworks are neural networks that generate the weights of another network conditioned on an input context vector. Crucially, they have already proven effective at resolving conflicting gradients in multi-task RL (Navon et al., 2020) and mitigating catastrophic forgetting in continual learning (von Oswald et al., 2020). In HyperMARL, they adaptively generate specialised or homogeneous policies and decouple observation- from agent-conditioned gradients (Section 4.4), a mechanism that empirically reduces gradient variance and improves policy updates (Sarafian et al., 2021). We further show that HyperMARL attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Section 6), suggesting it helps mitigate cross-agent gradient interference in FuPS.
@@
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation from the Vectorized Multi-Agent Simulator (VMAS) (Bettini et al., 2022), Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021), and SMAX (Rutherford et al., 2024)—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo (Bettini et al., 2024) and HAPPO (Zhong et al., 2024). Furthermore, HyperMARL achieves NoPS-level behavioural diversity while using a shared architecture – a challenge for conventional FuPS methods.
@@
  • We propose HyperMARL (Section 4), a method that leverages agent-conditioned hypernetworks to adaptively learn diverse, homogeneous or mixed behaviours, without altering the learning objective or requiring preset diversity levels.
  • We empirically evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo and HAPPO, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6). Additionally, HyperMARL scales efficiently by maintaining a near-constant number of parameters as the number of agents increases (Section C.2).
@@ 2 Background
We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP (Oliehoek & Amato, 2016). A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e. $O(\mathbf{o},\mathbf{a},s)=\Pr(\mathbf{o_{t}}|s_{t},\mathbf{a_{t-1}})$, $T:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the state transition function i.e. $T\left(s,\mathbf{a},s^{\prime}\right)=\Pr\left(s_{t+1}\mid s_{t},\mathbf{a}_{t}\right)$ and $\gamma$ is the discount factor.

### v3

@@ Abstract
Adaptability to specialised or homogeneous behaviours is critical in cooperative multi-agent reinforcement learning (MARL). Parameter sharing (PS) techniques, common for efficient adaptation, often limit behavioural diversity due to cross-agent gradient interference, which we show can be exacerbated by the coupling of observations and agent IDs. Current remedies typically add complexity through altered objectives, manual preset diversity levels, or sequential updates. We ask: can shared policies adapt without these complexities? We propose HyperMARL, a PS approach using hypernetworks for dynamic agent-specific parameters, without altering the RL objective or requiring preset diversity levels. HyperMARL’s explicit decoupling of observation- and agent-conditioned gradients empirically reduces policy gradient variance, facilitates shared-policy adaptation (including specialisation), and helps mitigate cross-agent interference. Across diverse MARL benchmarks (up to 20 agents), requiring homogeneous, heterogeneous, or mixed behaviours, HyperMARL achieves competitive performance against key baselines – fully shared, non-parameter sharing, and three diversity-promoting methods – while preserving behavioural diversity comparable to non-parameter sharing. These findings establish HyperMARL as a versatile approach for adaptive MARL. The code is publicly available at https://github.com/KaleabTessera/HyperMARL.
@@ 1 Introduction
Specialist and generalist behaviours are critical to collective intelligence, enhancing performance and adaptability in both natural and artificial systems [47, 38, 39, 22, 45]. In Multi-Agent Reinforcement Learning (MARL), this translates to a critical need for policies that can flexibly exhibit specialised, homogeneous, or mixed behaviours to meet diverse task demands [25, 6, 2].
@@
Optimal MARL performance thus hinges on being able to represent the required behaviours. While No Parameter Sharing (NoPS) [28] enables specialisation by using distinct per-agent networks, it suffers from significant computational overhead and sample inefficiency [10]. Conversely, Full Parameter Sharing (FuPS) [40, 17, 14], which trains a single shared network, improves efficiency but typically struggles to foster the behavioural diversity necessary for many complex tasks [23, 16, 25].
@@
Balancing FuPS efficiency with the capacity for diverse behaviours therefore remains a central open problem in MARL. Prior works have explored intrinsic-rewards [25, 21], role-based allocations [43, 44], specialised architectures [23, 26, 6], or sequential updates [49]. However, these methods often alter the learning objective, require prior knowledge of optimal diversity levels, involve delicate hyperparameter tuning, or necessitate maintaining agent-specific parameters and sequential updates (see Section 7).
@@
These challenges lead us to ask: Can we design a shared MARL architecture that flexibly supports both specialised and homogeneous behaviours—without altered learning objectives, manual preset diversity levels, or sequential updates? A key difficulty with FuPS, particularly for diverse behaviours, was hypothesised to be gradient interference among agents, whereby their updates negatively impact each other’s learning [10, 49]. We not only empirically validate this hypothesis but also demonstrate a critical insight: this conflict is significantly exacerbated by the common practice of coupling observations with agent IDs within a shared network (Fig. 1 for coupling, Sec. 3.2 for results).
@@
To counteract the challenges posed by this coupling, we propose HyperMARL, a novel parameter sharing method that explicitly decouples observation- and agent-conditioned gradients (Section 4.3) through hypernetworks [18]. Specifically, agent-conditioned hypernetworks, which condition on an agent identifier (or embedding), and generate an agent’s actor and critic weights on the fly (Fig. 1). Hypernetworks have already proved effective for resolving gradient conflicts in multi-task RL [30] and for mitigating catastrophic forgetting in continual learning [42], our work establishes their effectiveness in MARL. We further show that HyperMARL empirically attains lower policy gradient variance than FuPS and that this decoupling is critical for specialisation (Sec. 5.26), suggesting it helps mitigate cross-agent gradient interference in shared architectures.
@@
We validate HyperMARL on diverse MARL benchmarks—including Dispersion and Navigation (VMAS) [5], Multi-Agent MuJoCo (MAMuJoCo) [34], and SMAX [36]—across environments with two to twenty agents that require homogeneous, heterogeneous, or mixed behaviours. HyperMARL consistently matches or outperforms NoPS, FuPS, and diversity-promoting methods such as DiCo [6],HAPPO [49] and Kaleidoscope [26], while achieving NoPS-level behavioural diversity while using a shared architecture.
@@
  • Motivated by evidence that coupling agent IDs with observations can be detrimental, we propose HyperMARL (Sec. 4), a method that leverages agent-conditioned hypernetworks to decouple observation- and agent-conditioned gradients. HyperMARL adaptively learns diverse, homogeneous, or mixed behaviours without altering the RL objective or requiring preset diversity levels.
  • We thoroughly evaluate HyperMARL (Section 5) across diverse MARL environments with up to 20 agents. HyperMARL consistently achieves competitive mean episode returns compared to NoPS, FuPS, and diversity-focused methods such as DiCo, HAPPO and Kaleidoscope, while achieving NoPS-level behavioural diversity.
  • We demonstrate that HyperMARL empirically reduces policy gradient variance compared to FuPS and facilitates specialisation, suggesting the importance of gradient decoupling for mitigating cross-agent interference (Section 5.2; Section 6).
@@ 2 Background
We formulate the fully cooperative multi-agent systems addressed in our work as a Dec-POMDP [31]. A Dec-POMDP is a tuple, $\langle\mathbb{I},\mathbb{S},\{\mathbb{A}^{i}\}_{i\in\mathbb{I}},R,\{\mathbb{O}^{i}\}_{i\in\mathbb{I}},O,T,\rho_{0},\gamma\rangle$, where $\mathbb{I}$ is the set of agents of size $n=|\mathbb{I}|$, $\mathbb{S}$ is the set of global states with an initial state distribution $\rho_{0}$, $\mathbb{A}^{i}$ is the action space for agent $i$ where $\mathbb{A}=\times_{i}\mathbb{A}^{i}$ is the joint action space, $R:\mathbb{S}\times\mathbb{A}\to\mathbb{R}$ is the shared reward function, $\mathbb{O}^{i}$ is the observation space for agent $i$ with the joint observation space $\mathbb{O}=\times_{i}\mathbb{O}^{i}$, $O:\mathbb{O}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the probability of joint observation $\mathbf{o}\in\mathbb{O}$, i.e. $O(\mathbf{o},\mathbf{a},s)=\Pr(\mathbf{o_{t}}|s_{t},\mathbf{a_{t-1}})$, $T:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\to[0,1]$ is the state transition function i.e. $T\left(s,\mathbf{a},s^{\prime}\right)=\Pr\left(s_{t+1}\mid s_{t},\mathbf{a}_{t}\right)$ and $\gamma$ is the discount factor.